OpenStudy (anonymous):

Essay question: show all work. Suzie's flower shop estimates that supplies costs for one month will be given by the polynomial 2x^2 +14x -9. She estimates that utilities costs for one month is given by 12x-3. Find and simplify an expression that would estimate Suzie's total costs for one month. I would like to know the expression, I probably could work it out, I suck at word problems, so its hard for me to get an expression out of words. Thank you for your help!

OpenStudy (anonymous):

Two expression are both cost so to get total we get add the expression

OpenStudy (anonymous):

Supply Cost = 2x^2+14x-9 Utilities cost = 12x-3 total cost = supply cost + utilities cost = 2x^2+14x-9+12x-3

OpenStudy (anonymous):

ok so: (2x^2+14x-9) +(12x-3)

OpenStudy (anonymous):

right. you can add up the like terms. 14x+12x = 26x

OpenStudy (anonymous):

So I have: =(2x^2+14x-9) + (12x-3) =2x^2+14x-9+12x-3 =2x^2+(14+12)x + (-9-3) =2x^2+26x-12

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you can go one step further and factor out 2 to get 2(x^2+13x-6)

OpenStudy (anonymous):

I don't know if i am suppose to. How would I know? I am doing polynomials.

OpenStudy (anonymous):

it doesn't matter. The meaning of the polynomial has not changed. you can do it if you want or leave it alone.

OpenStudy (anonymous):

ok, thanks again.

OpenStudy (anonymous):

ok, thanks again.